OpenStudy (anonymous):

solving this quadratic equation by factoring (x+1)(2x-3)=8(2x-3)

OpenStudy (anonymous):

is this the question or you stuck here???

OpenStudy (anonymous):

im stuck on this question...but the topic is called "solving quadratic equations by factoring

OpenStudy (anonymous):

when you will see the discriminant, it is less than 0. so it can not be factorize| it has complex solutions.

OpenStudy (anonymous):

so than wat do I do? :(

OpenStudy (anonymous):

you can find the answer by this formula

OpenStudy (anonymous):

have you studied complex numbers???

OpenStudy (anonymous):

nope :(

OpenStudy (anonymous):

oh i made a mistake the answer is x=3/2 and 7

OpenStudy (anonymous):

(x+1)(2x-3)=8(2x-3) 2x^2 -3x+2x-3=16x-24 2x^2 -x-3-16x+24=0 2x^2 -17x+21=0 2x^2-14x-3x+21=0 2x(x-7)-3(x-7)+0 (2x-3)(x-7)=0 x=(3/2) and x=7

OpenStudy (anonymous):

thankxxxxx

OpenStudy (anonymous):

@RebeL_NYC :-)

OpenStudy (anonymous):

how bout xsquared=81

OpenStudy (anonymous):

+9 ans -9